OpenStudy (anonymous):

find the derivative of x√(1+y)+y√(1+x)

OpenStudy (anonymous):

you must specify the variable with respect to which u need to find the derivative...

OpenStudy (anonymous):

that's the problem. it is not specified in the question

OpenStudy (anonymous):

do you have an equation or just an expression? shouldn't this be set equal to something?

OpenStudy (anonymous):

whatever i've typed above is the question. there's nothing more to it.

OpenStudy (anonymous):

lets assume that it is an equation and solve. then the answer should be \[-(\sqrt{1+y}+(0.5*y/\sqrt{1+x}))/(\sqrt{1+x}+(0.5*x/\sqrt{1+y}))\]